when i was two, my brother was half my age
now i am 100. what is my brothers age

Found 3 solutions by ptaylor, Alan3354, sachi:
Answer by ptaylor(2198)

Let x=your brother's age
Now the problem tells us that 98 years ago, your were 2 years old and your brother was (1/2)*2=1 year old
So our equation to solve is:
x-98=1 add 98 to each side
x=99 years old----your brother's age
Hope this helps---ptaylor

Answer by Alan3354(69443)

when i was two, my brother was half my age
You're 1 year older.
---------
now i am 100. what is my brothers age
You're still 1 year older.

Answer by sachi(548)

when i was two, my brother was half my age
now i am 100. what is my brothers age
ans:
when you are two your brother was half of you means 1 year
that means your brother is 1 year junior to you
so now when you are 100 your brother must be 100-1=99 years old ans